A hockey puck leaves the blade of a hockey stick, bounces off a wall, and travels in a new direction.
a) Write and equation that models the path of the ordered pairs of (-4,0) and (0,8).
b) Write and equation that models the path of the ordered pairs of (0,8) and (-4,16).
c) Does the path of the puck form a right angle? Explain.

Answer by stanbon(75887)

a) Write and equation that models the path of the ordered pairs of (-4,0) and (0,8).
slope = (8-0)/(0--4) = 2
intercept: Since f(0) = 8,8 is the intercept
Equation:
y = 2x + 8
-----------------------------------------

b) Write and equation that models the path of the ordered pairs of (0,8) and (-4,16).
slope = (16-8)/(-4-0) = 8/-4 = -2
intercept: since f(0) = 8, the intercept is .
Equation:
y = -2x+8
-----------------------------------------
c) Does the path of the puck form a right angle? Explain.
No, because the product of the slopes is not -1.
====================================
Cheers,
Stan H.